# Fast SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

### v1

@@ 6.3 Proton transport in heterogeneous medium
To assess the capability of the proposed model to handle material heterogeneities and associated range shifts, a second test was conducted using a cubic phantom containing a 2-cm-thick bone insert positioned at the beam entrance. All other parameters were kept identical to those of the homogeneous water case described in the previous section. Figure 8 presents the 1D comparisons for 100 and 150 MeV monoenergetic proton beams. The integrated depth–dose curves show excellent agreement between the two models, with calculated range differences below 0.1 mm for 100 MeV and 0.4 mm for 150 MeV. Along the central axis, the dose deviations within the bone region remain below 5%. Interestingly, this region exhibits a slight underestimation by the SDE model, in contrast to the homogeneous case, where an overall dose overestimation was observed. As illustrated by the lateral profiles in Figure 9, this underestimation is confined to the bone region of the phantom. Beyond the bone interface, where the medium returns to water, both models regain strong agreement at mid-depths, while the SDE again shows an overestimation near the Bragg peak.
(a) Integral depth-dose comparison.
(b) Central-axis depth-dose distribution comparison.
@@
Figure 8: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a heterogeneous phantom with a 2-cm bone layer. Lower subplots show pointwise percentage dose differences.
(a) Profiles using 100 MeV proton beam at different depths.
(b) Profiles using 150 MeV proton beam at different depths.
@@
Figure 9: Lateral profile comparison between SDE (marker points) and Geant4 (solid lines) for two proton energies in a heterogeneous phantom. The profile at 1 cm corresponds to the bone region.
To further quantify agreement, a gamma analysis was performed for both beam energies in the heterogeneous configuration, as shown in Figure 10 and summarised in Table 2. For clinically relevant criteria (3%/2 mm, 10% dose threshold), pass rates exceeded 99%, indicating strong consistency between SDE and Geant4 dose distributions. Even when applying more strict criteria (2%/1 mm, 10% dose threshold), the pass rates remained above 97%. Spatially, the bone region shows gamma indices close to but below 1 for the 3%/2 mm criterion, indicating that the agreement between the SDE and Geant4 remains within acceptable limits even in high-density regions. However, for more stringent criteria, some voxels in the bone region begin to fail, suggesting that further refinement may be needed to improve accuracy in high-density media.
(a) Gamma distribution using 100 MeV proton beam.
(b) Gamma distribution using 150 MeV proton beam.
Figure 10: Central-axis gamma index matrices for 100 MeV and 150 MeV pristine proton beams in heterogeneous phantom.
Energy (MeV) DD (%) DTA (mm) Threshold (%) Norm. Method Pass rate (%)
100 3 2 10 global 99.53
3 2 10 local 99.53
3 2 1 local 99.41
2 1 1 local 97.52
150 3 2 10 global 99.62
3 2 10 local 99.38
3 2 1 local 99.50
2 1 1 local 98.86
Table 2: Gamma analysis pass rates in heterogeneous phantom using two monoenergetic proton energies.
6.4 Dose distribution in low density materials
The SDE underestimates the radial spread of dose in low-density media, such as air, due to the fact that it does not track ionisation electrons. Instead, the energy which would give rise to these electrons is deposited locally at the position of the primary proton. Ionisation electrons can travel a long distance when the density of the medium is low, resulting in a discrepancy in the radial spread of dose between the SDE output and a Geant4 benchmark.
We illustrate this effect by considering the same bone-water phantom as in Section 6.3 with a 5 cm air gap between the nozzle and the bone layer. Integral depth-dose curves (Figure 11(a)) show good overall agreement between the SDE and Geant4, although slight discrepancies are observed at shallow depths, with differences approaching 20%. Central-axis depth-dose curves (Figure 11(b)) indicate that dose differences exceed 20% within the air region. We confirmed this by changing the electron production cut in Geant4 from the default value of 1 mm to a very large value (100 cm), the effect of which is for Geant4 to neglect these electrons and instead deposit their dose locally at the position of their creation. This removed the discrepancy, producing results similar to the SDE output.
(a) Integral depth-dose comparison.
(b) Central-axis depth-dose distribution comparison.
@@ 6.4 Dose distribution in low density materials
Figure 11: 1D comparison between SDE and Geant4 for a 100 MeV beam in a heterogeneous phantom with a 2-cm bone layer and an additional 5 cm air gap at the beam entrance. Lower subplots show pointwise percentage dose differences.
(a) Lateral profile comparison between SDE (marker points) and Geant4 (solid lines) at different depths.
(b) Central-axis gamma index matrix.
Figure 12: Further comparisons performed for a 100 MeV beam in a heterogeneous phantom, adding a 5 cm air gap before the phantom. The profiles at 2.5, 6.0 and 11.5 cm correspond to air, bone and water regions.
Despite this discrepancy, the proton range agreement between the SDE and Geant4 remains within 0.2 mm, and lateral spread maintains good agreement in the denser materials, as seen in Figure 12(a), which demonstrates no change in bone or water regions. However, at the air depth (2.5 cm), the lateral profile shows overestimation at the beam centre, consistent with the observed dose difference. Furthermore, 3D gamma analysis retains pass rates above 99% under strict criteria (3%/2 mm, 1% dose threshold), and the air region does not fail the gamma test, as seen in Figure 12(b). Gamma analysis passes even in air because the region in which the SDE overestimates dose is very highly localised to the central voxel of the beam. Hence, a strict 2 mm comparison radius suffices to smooth out its effect.
These findings indicate that the SDE model underestimates dose in low-density materials, but clinical impact is limited since treatment-relevant doses are delivered to tissue rather than air gaps. For organs with significant air volumes, such as lungs, the resulting discrepancy could be relevant. Future work could address this by incorporating a mathematical description of electron diffusion in low-Z materials, rather than tracking individual electrons as in Geant4.
6.5 Simulation time comparison
[…]
7 Conclusions
This study presents the first comprehensive validation of the SDE model for proton dose calculation, benchmarked against Geant4 in both homogeneous and heterogeneous phantoms. The model reproduces depth-dose curves, lateral spread, and 3D gamma indices with deviations that are comparable to those observed among various Geant4 physics lists. Additionally, the SDE model demonstrates a considerable computational speed advantage, which becomes more pronounced at higher proton energies. The implementation is readily parallelisable, which will allow for further speed increase. Future work will focus on simplified modelling of secondary particle contributions and optimisation of the source code for clinical use. Overall, these findings identify the SDE approach as a promising, fast and accurate alternative to conventional Monte Carlo simulations in proton therapy.
@@ A.2 Integration of (9) for Hydrogen elastic scattering
The authors acknowledge support from the EPSRC grant Mathematical Theory of Radiation Transport: Nuclear Technology Frontiers (MaThRad), EP/W026899/2. Moreover, we would like to thank colleagues from the UCLH proton beam facility (in particular Colin Baker and Sarah Osman) as well Ana Lourenço from the National Physical Laboratory for their guidance.

### v2

@@ 3.2 Longitudinally heterogeneous slab phantom
To assess the capability of the proposed model to handle multiple material interfaces and associated range shifts, a second test was conducted adding a bone slab and a low-density compressed lung slab to the water phantom. All other parameters were kept identical to those of the homogeneous water case described in the previous section. Figure 6 presents the 1D comparisons for 100 and 150 MeV monoenergetic proton beams. The integrated depth–dose curves show excellent agreement between the two models, with calculated range differences below 0.2 mm for 100 MeV and 0.4 mm for 150 MeV.
@@
Figure 6: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a water phantom with a 1-cm bone layer and a 2-cm low density lung layer. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
The lateral dose profiles for both proton energies in the slab phantom are shown in Figure 7, whereas the 2D dose distributions and voxel-wise relative difference maps for this scenario are shown in Figure 8. Overall, the behaviour is consistent with that observed in the homogeneous phantom, with the SDE model reproducing the lateral profile shapes and widths across all material layers. This indicates that the dominant beam broadening is well described throughout the heterogeneous geometry. Profiles are shown at representative depths within each slab layer, as well as at the depths corresponding to the Bragg peak positions. For the 100 MeV beam, the largest discrepancies are concentrated around the Bragg peak, as also observed in the homogeneous case. In the slab configuration, this difference is more noticeable due to a sharper Bragg peak at lower energy and the presence of multiple material interfaces, which make the peak height more sensitive to small differences in scattering and energy-deposition modelling, even when the proton range is well matched.
@@
Figure 7: Lateral profile comparison between SDE (marker points) and Geant4 (solid lines) for two proton energies in a longitudinally heterogeneous phantom (a) Using a 100 MeV beam, (b) using a 150 MeV beam. The profiles at 2.5 and 4 cm correspond to the bone and lung regions.
Figure 8: 2D dose distribution comparison between the SDE model and Geant4 in a slab phantom for monoenergetic proton beams. (a) Central-axis dose distribution and (b) corresponding relative dose difference map for a 100 MeV beam, normalised to the 99th percentile (D99) of the Geant4 reference dose. (c,d) Same quantities for a 150 MeV beam.
For the lower-energy beam, milder differences are observed in the low-density lung region between depths of 3 and 5 cm. In low-density media, energy deposition is more strongly influenced by how energy is redistributed away from the proton track. Geant4 explicitly transports secondary electrons, leading to a slightly broader lateral redistribution of dose, while the SDE model deposits energy more locally. This results in small but spatially coherent differences in the lung region.
For the higher-energy beam, differences in the lung region are less apparent, while the main discrepancies occur within the bone layer between depths of 2 and 3 cm. At higher energies, transport in low-density media becomes less sensitive to secondary-electron effects, whereas the increased stopping power and sharp density change in bone have a greater influence on the dose distribution. As a result, differences between the two models are more visible in the high-density layer.
To further quantify agreement, gamma analysis was performed for both beam energies in the slab configuration, yielding pass rates of 95.1% for the 100 MeV beam and 98.6% for the 150 MeV beam using a 2%/0.5 mm criterion with a 1% dose threshold. In terms of computational performance, the SDE model was between 2.6 and 3 times faster than single-threaded Geant4 for this geometry.
3.3 Laterally heterogeneous composite phantom
Having established good agreement between the SDE model and Geant4 for the homogeneous and slab phantom configurations, the laterally heterogeneous insert phantom provides a more stringent test of the model’s response to localised lateral heterogeneities, which are known to challenge reduced proton transport models (Schaffner et al., 1999). Figure 9 shows the integrated and central-axis dose distributions, which remain in close agreement and within the same percentage difference ranges observed for the previous test cases.
@@ 3.3 Laterally heterogeneous composite phantom
Figure 9: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a heterogeneous phantom with a 2-cm off-axis bone insert. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
Both models reproduce the two expected high-dose peaks arising from the range shift experienced by the upper half of the beam. The agreement in proton range for the more prominent peak is within 0.2 mm for the 100 MeV case and within 0.5 mm for the 150 MeV case, with a comparable level of agreement retained for the secondary peak.
Lateral dose profiles are shown in Figure 10 and illustrate the influence of the bone insert at greater depths. The two shallower profiles correspond to the entrance water region and the bone insert location, while the deeper profiles correspond to the two Bragg peak positions formed by the upper and lower portions of the beam. While the overall lateral extent of the dose distributions is comparable between the SDE model and Geant4, small differences are observed in the central region of the profiles. For the 100 MeV beam, a reduced dose is again observed for one of the Bragg peaks, whereas the secondary peak shows closer agreement, consistent with its broader spatial distribution. These differences reflect variations in the radial redistribution of deposited energy in the presence of localised heterogeneities, rather than changes in the overall lateral spread of the beam.
Figure 10: Lateral dose profile comparison between the SDE model (marker points) and Geant4 (solid lines) for two proton energies (a) 100 MeV and (b) 150 MeV in a laterally heterogeneous phantom at different depths. The profile at 4 cm corresponds to the bone insert region. The profiles at the two largest depths correspond to the locations of the high-dose peaks observed in the 1D depth–dose distributions.
The 2D dose distributions and voxel-wise relative difference maps for this configuration are shown in Figure 11. The effects of lateral heterogeneities are well captured by the SDE model, with relative differences remaining within the same limits observed for the homogeneous and slab phantoms. This indicates that the magnitude and spatial localisation of the model deviations are stable across the different phantom geometries considered.
Gamma analysis for the insert phantom yielded pass rates of 97% for the 100 MeV beam and 98.5% for the 150 MeV beam, using the same stringent criteria applied in the previous scenarios. In terms of computational performance, the SDE model remains between 2.5 and 2.9 times faster than single-threaded Geant4 for this configuration.
Figure 11: 2D dose distribution comparison between the SDE model and Geant4 in a water phantom with an off-axis bone insert for monoenergetic proton beams. (a) Central-axis dose distribution and (b) corresponding relative dose difference map for a 100 MeV beam, normalised to the 99th percentile (D99) of the Geant4 reference dose. (c,d) Same quantities for a 150 MeV beam.
4 Discussion and Conclusion
This study presents the first systematic validation of the SDE-based proton transport model against Geant4 across homogeneous and heterogeneous phantom geometries. The results demonstrate that the SDE model reproduces the principal dosimetric features predicted by Geant4 with consistent accuracy. Proton range agreement remained within 0.2 mm for 100 MeV beams and within 0.6 mm for 150 MeV beams across all test cases, and gamma pass rates exceeded 95% under stringent 2%/0.5 mm criteria with a 1% dose threshold—well within clinically acceptable bounds. Importantly, the observed differences between the two models were spatially stable and well characterised: discrepancies were concentrated around steep dose gradients and material interfaces, while entrance, plateau, and lateral regions showed close agreement. Comparison against multiple Geant4 physics lists further confirmed that the deviations introduced by the SDE model lie within the range of inter-list variability typical of Geant4 itself, reinforcing the physical consistency of the approach. The computational performance of the SDE model was also stable across all configurations, achieving speed-ups of approximately 2.5–3× relative to single-threaded Geant4, with the greatest gains observed at higher beam energies. It is important to frame the claim being made here precisely: the SDE formulation offers a genuinely novel application of stochastic differential equations to charged particle transport, and— independent of any hardware acceleration—it can simulate proton transport faster than conventional Monte Carlo codes for comparable dosimetric accuracy. While GPU-accelerated Monte Carlo engines are increasingly being adopted in clinical practice, a factor of approximately three in per-proton computational cost represents a meaningful qualitative improvement. Such gains could enable longer optimisation runs during treatment planning on more modest computing infrastructure that not all clinics have the ability to realise.
The simplifications adopted in the current model—notably the absence of secondary particle transport and the use of a fixed energy absorption threshold—were shown to be appropriate for producing clinically relevant outputs in the geometries and energy ranges considered. The largest resulting effects were a steeper distal fall-off and the absence of low-level dose contributions at remote locations, both of which are expected and well understood. Regarding model parameterisation, the central-axis depth–dose curves exhibited sensitivity to the angular cutoffs used for Rutherford and backscatter events, which are currently set as fixed values. Investigating energy-dependent cutoffs in future work may reduce the number of free parameters and improve robustness across a wider range of clinical scenarios. Beyond the single-threaded results reported here, the SDE framework is naturally amenable to further acceleration. Each sequential proton track evolves independently, making the model straightforward to parallelise across multiple CPU cores—a strategy routinely used to accelerate Monte Carlo simulations—or to port to GPU architectures. Aside from relatively rare inelastic proton-nucleus collision which can cause macroscopic losses in proton energy, each SDE step involves identical arithmetic operations across all protons, making the method well-suited to SIMD (Single Instruction, Multiple Data) execution. This suggests substantial additional speed improvements are achievable without fundamental algorithmic modifications, and positions the SDE framework as a natural candidate for GPU acceleration.
Several avenues for future development follow from the present work. Incorporating simplified models for secondary particle contributions, particularly neutrons and gamma rays, would improve agreement in the distal tail and at remote locations. Extending the validation to patient-derived CT geometries and spread-out Bragg peaks would provide a more complete assessment of clinical applicability. Such geometries will introduce additional challenges, including voxelised material assignment and the presence of air cavities with sharp density contrasts. The phantom-based validation presented here establishes the necessary foundation by demonstrating accuracy across controlled homogeneous and heterogeneous configurations before confronting these complexities. Beyond computational speed, the continuous and differentiable structure of the SDE formulation opens possibilities that are fundamentally difficult to achieve with discrete Monte Carlo sampling. The well-characterised probabilistic structure of the SDE provides a natural framework for propagating input uncertainties (such as stopping power ratios or patient positioning errors) directly through the transport model, complementing conventional scenario-based robustness evaluation strategies, e.g. optimising the worst-case scenario. Realising these capabilities in practice will require further methodology, but the mathematical foundation provided by the SDE framework makes such extensions feasible. Source code optimisation and integration with treatment planning workflows are also natural next steps. Overall, these findings establish the SDE approach as a promising, fast, and accurate complement to conventional Monte Carlo methods for proton therapy dose calculation.
@@ A.2 Integration of (9) for hydrogen elastic scattering
The authors acknowledge support from the EPSRC grant Mathematical Theory of Radiation Transport: Nuclear Technology Frontiers (MaThRad), EP/W026899/2. Moreover, we would like to thank colleagues from the UCLH proton beam facility (particularly Colin Baker and Sarah Osman), from the University of Cambridge (particularly Theophile Bonnet, Valeria Raffuzzi and Esmae Woods), as well Ana Lourenço from the National Physical Laboratory for help and guidance.